# Key varieties for prime $\mathbb{Q}$-Fano threefolds defined by Freudenthal triple systems

### v2

@@ 7.4. The case $A_{3}$ and $A_{4}$ are nonzero constants
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A},A_{3}\!A_{4}}^{8}$ is presented in the format of the equations of $\mathfrak{S}_{\mathbb{A}}^{8}$ by setting
$x={}^{t}\!\left(\begin{array}[]{ccc}A_{41}&\theta_{2}&\theta_{3}-2/3\lambda_{13}A_{41}+1/3\lambda_{24}\theta_{2}\end{array}\right),$
$y={}^{t}\!\left(\begin{array}[]{ccc}\theta_{1}&A_{23}&\theta_{4}+1/3\lambda_{13}\theta_{1}-2/3\lambda_{24}A_{23}\end{array}\right),$
@@
The affine variety $\mathfrak{Cl}_{\mathbb{A},A_{3}\!A_{4}}^{8}$ is isomorphic to $\mathfrak{S}_{\mathbb{A}}^{8}$. Moreover, the isomorphism is weighted homogeneous with the corresponding weights of coordinates of the equations as in Proposition 5.3.
@@
We consider the weighted projectivization $\mathfrak{S}_{\mathbb{P}}^{7}$ of $\mathfrak{S}_{\mathbb{A}}^{8}$ in $\mathbb{P}(1^{8},2^{4})$ by putting the weights of coordinates as follows:
@@
By [CD], this produces a prime $\mathbb{Q}$-Fano 3-fold $X$ of No.20652 as its weighted complete intersection by one hypersurface of weight 2 and three hypersurfaces of weight 1. It is easy to verify that the the set of three $1/2\,(1,1,1)$-singularities of $X$ is the intersection between the twisted cubic curve $\gamma:=\{U=O,D=0,\wedge^{2}V=\bm{o}\}$ and $X$.
@@
A prime $\mathbb{Q}$-Fano $3$-fold $X$ constructed in Example 7.9 belongs to the class No.4.1 of [Tak1] for any $1/2\,(1,1,1)$-singularity of $X$.
@@
We choose a $1/2\,(1,1,1)$-singularity $\mathsf{p}$ and construct the Sarkisov link starting from the blow-up at $\mathsf{p}$. By [Tak1, Tables 4 and 5], we can distinguish between prime $\mathbb{Q}$-Fano threefolds of No.4.1 and No.5.4 by the dimension of forms of weight 2 vanishing at $\mathsf{p}$ with weighted multiplicity $\geq$ $3$; the dimension is $3$ for No.4.1, and is 4 for No.5.4, which can be verified by looking at the rational map $X\dashrightarrow X^{\prime}$ in the Sarkisov link (note that this map is defined by the linear system $|-2K_{X}-3\mathsf{p}|)$. By the group action on $\mathfrak{S}_{\mathbb{A}}^{8}$ described in Subsection 5.2, we may assume that $\mathsf{p}$ is the $v_{0}$-point (the point whose coordinates except $v_{0}$ are zero). Then, by the equation (5.1), we see that 4 forms of weight 2, $u_{1}^{2},u_{1}u_{3},u_{3}^{2},v_{3}$ are all such forms and they are linearly independent on $\mathfrak{S}_{\mathbb{A}}^{8}$. We have only to show that they are still linearly independent on $X$. Assume the contrary. Then we may assume that the hypersurface $Q$ of weight 2 cutting $X$ from $\mathfrak{S}_{\mathbb{P}}^{7}$ is $\{av_{3}+bu_{1}^{2}+cu_{1}u_{3}+du_{3}^{2}=0\}$ with some $a,b,c,d\in\mathbb{C}$. Then $Q$ intersects the twisted cubic curve $\gamma$ only at the $v_{0}$-point, or contains $\gamma$. This is a contradiction since $Q$ must intersects $\gamma$ at three distinct points for $X$ to have three $1/2\,(1,1,1)$-singularities. ∎
@@
In [Tak6], we will construct a prime $\mathbb{Q}$-Fano threefold of No.5.4 via another key variety $\mathscr{H}_{\mathbb{A}}^{13}$.
7.5. The case $A_{1}$ and $A_{4}$, or $A_{2}$ and $A_{3}$ are nonzero constants
@@
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A},A_{1}\!A_{4}}^{8}$ is presented in the format of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ by the setting as in Proposition 7.5 with $A_{1}=-1$.
@@
By Proposition 7.2 and the assumption that $A_{1}=A_{4}=-1$, we see that the entries $\theta_{1}-1/3\lambda_{13}A_{23}$ and $-\theta_{4}+1/3\lambda_{24}A_{23}$ are weighted homogeneous. Moreover, we set
@@
Moreover, the isomorphism is weighted homogeneous with the corresponding weights of coordinates of these equations.
@@
7.6. The case $A_{1}$ and $A_{3}$ are nonzero constants– a subvariety of $\mathfrak{Z}_{\mathbb{A}}^{12}$
@@
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$ is presented in the format of the equations of $\mathfrak{Z}_{\mathbb{A}}^{12}$ by setting
$x={}^{t}\!\left(\begin{array}[]{ccc}\theta_{4}&\theta_{1}&A_{23}\end{array}\right),y={}^{t}\!\left(\begin{array}[]{ccc}\theta_{2}&A_{41}&\theta_{3}-\lambda_{13}A_{41}\end{array}\right),$
@@
of $\mathfrak{Z}_{\mathbb{A}}^{12}$. Moreover, the isomorphism is weighted homogeneous with the corresponding weights of coordinates of the equations as in Definition 6.1.
@@ 7.7. The case with the most number of nonzero constant coordinates
we see by a straightforward calculation that $\mathfrak{S}_{\mathbb{A}}^{8}\cap\{d_{3}=-1\}$ is isomorphic to the cone over $\mathfrak{S}_{\mathbb{A}}^{8}\cap\{d_{2}=0,\,d_{3}=-1\}.$ The isomorphism is weighted homogeneous with the corresponding weights of coordinates since $D_{0},D_{1},U_{1},U_{2}$ are weighted homogeneous if $A_{1}=A_{3}=A_{4}=-1$ by Proposition 7.2.
@@
Let $\rho_{\mathfrak{B}}:\widehat{\mathfrak{B}}\to\{d_{2}=0,d_{3}=-1\}$ be the base change of $\rho_{\mathfrak{S}}$ as in Subsection 5.3 by the inclusion map $\{d_{2}=0,d_{3}=-1\}\hookrightarrow\mathbb{A}_{D}^{4}$. By Proposition 5.8, we have immediately the following noting that the cone over the tangential scroll (5.2) of the twisted cubic restricts to the affine cuspidal cubic curve $\{d_{2}=0,d_{3}=-1,d_{0}^{2}=4d_{1}^{3}\}$:
@@ 7.7 $A_{1}$, $A_{3}$ and $A_{4}$ $35$
In the paper [P3], Papadakis constructs affine varieties via type ${\rm II}_{2}$ unprojection and proves they are Gorenstein (see [P3, Thm.2.15]). Moreover, in [P4], he provides explicit descriptions of parts of their equations, and a full description of the equation in one particular case (see [P4, Sect.4]), for which we denote by $\mathfrak{P}_{\mathbb{A}}^{23}$ the affine variety he constructs. The affine variety $\mathfrak{P}_{\mathbb{A}}^{23}$ is a subvariety of codimension 4 in the affine $27$-space with the coordinates implemented in the following format:
@@
    In [P4, Sec.5], two candidates of prime $\mathbb{Q}$-Fano 3-folds of anti-canonical codimension 4 with type ${\rm II}_{1}$ projections are constructed from $\mathfrak{P}_{\mathbb{A}}^{23}$. In her phd thesis submitted to Warwick University [Tay], Taylor constructs more such candidates. The remaining problem is to show their examples have Picard number 1.

### v3

@@ 7.4. The case where $A_{3}$ and $A_{4}$ are nonzero constants
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A},A_{3}\!A_{4}}^{8}$ are presented in the format of the equations of $\mathfrak{S}_{\mathbb{A}}^{8}$ by setting
[…]
@@
The affine variety $\mathfrak{Cl}_{\mathbb{A},A_{3}\!A_{4}}^{8}$ is isomorphic to $\mathfrak{S}_{\mathbb{A}}^{8}$. Moreover, the isomorphism is weighted homogeneous with the corresponding weights on the coordinates of the equations as in Proposition 5.3.
@@
We consider the weighted projectivization $\mathfrak{S}_{\mathbb{P}}^{7}$ of $\mathfrak{S}_{\mathbb{A}}^{8}$ in $\mathbb{P}(1^{8},2^{4})$ by putting the weights on the coordinates as follows:
@@
By [CD], this produces a prime $\mathbb{Q}$-Fano 3-fold $X$ of No.20652 as its weighted complete intersection by one hypersurface of weight 2 and three hypersurfaces of weight 1. It is easy to verify that the the set of three 1/2(1,1,1)1/2\,(1,1,1)-singularities of $X$ is the intersection between the twisted cubic curve $\gamma:=\{U=O,D=0,\wedge^{2}V=\bm{o}\}$ and $X$.
@@
A prime $\mathbb{Q}$-Fano $3$-fold $X$ constructed in Example 7.9 belongs to the class No.4.1 of [Tak1] for any 1/2(1,1,1)1/2\,(1,1,1)-singularity of $X$.
@@
We choose a 1/2(1,1,1)1/2\,(1,1,1)-singularity $\mathsf{p}$ and construct the Sarkisov link starting from the blow-up at $\mathsf{p}$. By [Tak1, Tables 4 and 5], we can distinguish between prime $\mathbb{Q}$-Fano threefolds of No.4.1 and No.5.4 by the dimension of forms of weight 2 vanishing at $\mathsf{p}$ with weighted multiplicity $\geq$ 33; the dimension is $3$ for No.4.1, and is 4 for No.5.4, which can be verified by looking at the rational map $X\dashrightarrow X^{\prime}$ in the Sarkisov link (note that this map is defined by the linear system $|-2K_{X}-3\mathsf{p}|)$. By the group action on $\mathfrak{S}_{\mathbb{A}}^{8}$ described in Subsection 5.2, we may assume that $\mathsf{p}$ is the $v_{0}$-point (the point whose coordinates except $v_{0}$ are zero). Then, by the equation (5.1), we see that 4 forms of weight 2, $u_{1}^{2},u_{1}u_{3},u_{3}^{2},v_{3}$ are all such forms and they are linearly independent on $\mathfrak{S}_{\mathbb{A}}^{8}$. We have only to show that they are still linearly independent on $X$. Assume the contrary. Then we may assume that the hypersurface $Q$ of weight 2 cutting $X$ from $\mathfrak{S}_{\mathbb{P}}^{7}$ is $\{av_{3}+bu_{1}^{2}+cu_{1}u_{3}+du_{3}^{2}=0\}$ with some $a,b,c,d\in\mathbb{C}$. Then $Q$ intersects the twisted cubic curve $\gamma$ only at the $v_{0}$-point, or contains $\gamma$. This is a contradiction since $Q$ must intersects $\gamma$ at three distinct points for $X$ to have three 1/2(1,1,1)1/2\,(1,1,1)-singularities. ∎
@@
In [Tak4], we will construct a prime $\mathbb{Q}$-Fano threefold of No.5.4 via another key variety $\mathscr{H}_{\mathbb{A}}^{13}$.
7.5. The case where $A_{1}$ and $A_{4}$, or $A_{2}$ and $A_{3}$ are nonzero constants
@@
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A},A_{1}\!A_{4}}^{8}$ are presented in the format of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ by the setting as in Proposition 7.5 with $A_{1}=-1$.
@@
By Proposition 7.2 and the assumption that $A_{1}=A_{4}=-1$, we see that the entries $\theta_{1}-1/3\lambda_{13}A_{23}$ and $\theta_{4}+1/3\lambda_{24}A_{23}$ are weighted homogeneous. Moreover, we set
@@
Moreover, the isomorphism is weighted homogeneous with the corresponding weights on the coordinates of these equations.
@@
7.6. The case where $A_{1}$ and $A_{3}$ are nonzero constants– a subvariety of $\mathfrak{Z}_{\mathbb{A}}^{12}$
@@
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are presented in the format of the equations of $\mathfrak{Z}_{\mathbb{A}}^{12}$ by setting
[…]
@@
of $\mathfrak{Z}_{\mathbb{A}}^{12}$. Moreover, the isomorphism is weighted homogeneous with the corresponding weights on the coordinates of the equations as in Definition 6.1.
@@ 7.7. The case with the most number of nonzero constant coordinates
we see by a straightforward calculation that $\mathfrak{S}_{\mathbb{A}}^{8}\cap\{d_{3}=-1\}$ is isomorphic to the cone over $\mathfrak{S}_{\mathbb{A}}^{8}\cap\{d_{2}=0,\,d_{3}=-1\}.$ The isomorphism is weighted homogeneous with the corresponding weights on the coordinates since $D_{0},D_{1},U_{1},U_{2}$ are weighted homogeneous if $A_{1}=A_{3}=A_{4}=-1$ by Proposition 7.2.
@@
[…]
@@ 7.7 $A_{1}$, $A_{3}$ and $A_{4}$ $35$
In the paper [P3], Papadakis constructs affine varieties via type ${\rm II}_{2}$ unprojection and proves they are Gorenstein (see [P3, Thm.2.15]). Moreover, in [P4], he provides explicit descriptions of parts of their equations, and a full description of the equation in one particular case (see [P4, Sec.4]), for which we denote by $\mathfrak{P}_{\mathbb{A}}^{23}$ the affine variety he constructs. The affine variety $\mathfrak{P}_{\mathbb{A}}^{23}$ is a subvariety of codimension 4 in the affine $27$-space with the coordinates implemented in the following format:
@@
    In [P4, Sec.5], two candidates of prime $\mathbb{Q}$-Fano 3-folds of anti-canonical codimension 4 with type ${\rm II}_{1}$ projections are constructed from $\mathfrak{P}_{\mathbb{A}}^{23}$. In her phD thesis submitted to Warwick University [Tay], Taylor constructs more such candidates. The remaining problem is to show that their examples have Picard number 1.